# Accurate modeling of continuous-time SAT solvers in SPICE

### v1

@@ Abstract
Recently, there has been an increasing interest in employing dynamical systems as solvers of NP-complete problems. In this paper, we present accurate implementations of two continuous-time dynamical solvers, known in the literature as analog SAT and digital memcomputing, using advanced numerical integration algorithms of SPICE circuit simulators. For this purpose, we have developed Python scripts that convert Boolean satisfiability (SAT) problems into electronic circuits representing the analog SAT and digital memcomputing dynamical systems. Our Python scripts process conjunctive normal form (CNF) files and create netlists that can be directly imported into LTspice. We explore the SPICE implementations of analog SAT and digital memcomputing solvers by applying these to a selected set of problems and present some interesting and potentially useful findings related to digital memcomputing and analog SAT.
@@ I Introduction
The structure of this paper is as follows. Section II provides an overview of the problem, solvers, and information on the SPICE modeling approach. Examples of LTspice simulations are given in Sec. III. This is followed by a discussion in Sec. IV. The paper ends with an acknowledgment.
@@ II-B2 Digital memcomputing
The dynamical system in the digital memcomputing approach [12] is characterized by Eqs. (3)-(5) (the original notation is used). In Eqs. (3)-(5), $v_{n}$-s are continuous variables (similar to $s_{i}$-s in the analog SAT), $x_{s,m}$-s and $x_{l,m}$-s are the short and long memory variables, and $q_{j,m}$-s are the constants defining the clauses (same as $c_{mj}$-s in the analog SAT). The performance of digital memcomputing depends on the constants $\alpha$, $\beta$, $\gamma$, $\delta$, $\epsilon$, and $\zeta$ [12]. The intervals for $v_{n}$, $x_{s,m}$, and $x_{l,m}$ are provided in Table I. Moreover, Eq. (10) defines the clause function, $C_{m}(v_{i},v_{j},v_{k})$, used in Eqs. (4), (5), and (9). This function characterizes the state of the variable that most closely satisfies clause $m$. For more information, see Refs. [12, 7].
@@ III-A Main simulations
(a)Refer to caption  (b)Refer to caption
(c)Refer to caption  (d)Refer to caption (e) Refer to caption
(f) Refer to caption
@@
(a)Refer to caption  (b)Refer to caption

### v2

@@ Abstract
Recently, there has been an increasing interest in employing dynamical systems as solvers of NP-complete problems. In this paper, we present accurate implementations of two continuous-time dynamical solvers, known in the literature as analog SAT and digital memcomputing, using advanced numerical integration algorithms of SPICE circuit simulators. For this purpose, we have developed Python scripts that convert Boolean satisfiability (SAT) problems into electronic circuits representing the analog SAT and digital memcomputing dynamical systems. Our Python scripts process conjunctive normal form (CNF) files and create netlists that can be directly imported into LTspice. We explore the SPICE implementations of analog SAT and digital memcomputing solvers by applying these to a selected set of problems and present some interesting and potentially useful findings related to digital memcomputing and analog SAT. In this work, we also introduce networks of continuous-time solvers with potential applications extending beyond the solution of Boolean satisfiability problems.
@@ I Introduction
The structure of this paper is as follows. Section II provides an overview of the problem, solvers, and information on the SPICE modeling approach. Examples of LTspice simulations are given in Sec. III. These include the application of solvers to a variety of SAT problems, modification of solvers, and simulation of digital memcomputing networks. Sec. III is followed by a discussion in Sec. IV. The paper ends with an acknowledgment.
@@ II-B2 Digital memcomputing
The dynamical system in the digital memcomputing approach [12] is characterized by Eqs. (3)-(5) (the original notation is used). In Eqs. (3)-(5), $v_{n}$-s are continuous variables (similar to $s_{i}$-s in the analog SAT), $x_{s,m}$-s and $x_{l,m}$-s are the short and long memory variables, and $q_{j,m}$-s are the constants defining the clauses (same as $c_{mj}$-s in the analog SAT). The performance of digital memcomputing depends on the parameters $\alpha$, $\beta$, $\gamma$, $\delta$, $\epsilon$, and $\zeta$ [12]. The intervals for $v_{n}$, $x_{s,m}$, and $x_{l,m}$ are provided in Table I. Moreover, Eq. (10) defines the clause function, $C_{m}(v_{i},v_{j},v_{k})$, used in Eqs. (4), (5), and (9). This function characterizes the state of the variable that most closely satisfies clause $m$. For more information, see Refs. [12, 7].
@@ II-D Generation of instances
(a)Refer to caption  (b)Refer to caption
(c)Refer to caption  (d)Refer to caption (e) Refer to caption
(f) Refer to caption
@@ III Simulations
In this section, we present examples of SPICE simulations for both the analog SAT and digital memcomputing solvers. Unless specified otherwise, the digital memcomputing results were obtained using the following parameter values: $\alpha=5$, $\beta=20$, $\gamma=0.25$, $\delta=0.05$, $\epsilon=0.001$, and $\zeta=0.01$.
@@ III-A Main simulations
(a)Refer to caption
(b)Refer to caption
@@ III-C2 Analog SAT
III-D Networks of SAT solvers
(a)Refer to caption   (c)Refer to caption
(b)Refer to caption   (d)Refer to caption
Figure 5: Solvers A and B’s response to a square wave voltage (represented by $v_{1}$). For the sake of clarity, the applied voltage curves were shifted downward by $2$ V. See Sec. III-D for more details. These simulations were performed using $\alpha=0$.
It might be worthwhile to explore networks composed of continuous-time SAT solvers, where these solvers serve as nodes that interact via some of their variables or combinations of them. On the one hand, SAT solver networks can be utilized to solve SAT problems whose size exceeds the capacity of an individual SAT solver, a particularly crucial approach for physical implementations of these solvers that have finite sizes [27]. On the other hand, SAT solvers can be viewed as distant analogs to neurons, possessing complex functionalities determined by the SAT problems they solve. Therefore, one can anticipate that networks of SAT solvers could find applications in artificial neural network architectures or reservoir computing.
An especially useful feature of SPICE for this application is the subcircuit declaration. Fundamentally, the subcircuit functions as a collection of elements in SPICE that can be referred to in a manner similar to device models. Externally, the subcircuit is accessed through its external nodes. For SAT solvers, a potential approach is to define $P$ input nodes and $Q$ output nodes ($P+Q\leq N$), such that, in the digital memcomputing case, the input nodes simply set the values of $P$ voltage variables (their dynamics equations are disregarded), while the output nodes are used to output the values of some other $Q$ voltage variables. To monitor the solver’s state, the signal $V_{\textnormal{contrd}}(t)$ may be included in the output nodes. Of course, there are other choices available to connect the solvers.
(a)        Refer to caption           (c) Refer to caption
(b) Refer to caption
Figure 6: (a) Schematics of a simple network of two digital memcomputing solvers. (b) Number of unsatisfied clauses as a function of time when $v_{2}$ is used as the output node in both solvers. (c) Number of unsatisfied clauses as a function of time when $v_{3}$ is used as the output node in both solvers.
As an example, we considered two instances, say A and B, of the 3-regular 3-XORSAT problem with $N=10$. The digital memcomputing circuits for A and B were created using our Python script [33]. First, we simulated uncoupled solvers and found the following solutions (which are not necessarily unique):
[…]
Here, each digit from left to right represents $x_{1}$ through $x_{10}$. By modifying several lines in the netlists of solvers A and B, we created the subcircuits for subsequent simulations. The sets $\{v_{1}\}$, $\{v_{i},V_{\textnormal{contrd}}(t)\}$ with $i=2$ or $3$ were selected as input and output nodes, respectively.
Next, we studied how individual solvers react to an applied square wave voltage. To achieve this, a square wave voltage source ($-1$ V to $+1$ V) was connected to the solver’s input node. We have observed that when the applied voltage is at logic zero and there is sufficient time, both solvers can successfully identify the solution. When the applied voltage is at logic one, the solvers exhibit some interesting dynamics, which can be described as deterministic for solver A and chaotic for solver B. These observations are evident in Fig. 5. A notable characteristic associated with solver A is the period doubling behavior depicted in Fig. 5(a), which vanishes at a longer period, as demonstrated in Fig. 5(b).
Finally, the solvers were connected in a network as shown in Fig. 6(a). Two cases were considered: in the first case, $v_{2}$ was used as the output node in both solvers, while in the second case, $v_{3}$ was employed for the same purpose. The networks were simulated for 300 time units of the circuit dynamics. Fig. 6(b) demonstrates that selecting $v_{2}$ as the output node results in no solution for the whole network, whereas using $v_{3}$ as the output node successfully yields a solution. These observations align with the solutions provided by Eqs. (15) and (16) and can be explained by them.
@@ IV Discussion
Lastly, we have demonstrated that it is quite straightforward to simulate networks of continuous-time solvers in SPICE. It is expected that such networks may have potential applications extending beyond the solution of Boolean satisfiability problems.